# Robot Cell Modeling via Exploratory Robot Motions

### v1

@@ Abstract
Index Terms:
Collision Avoidance, Physical Human-Robot Interaction, Software Tools for Robot Programming.
@@ I Introduction
Figure 1: Swept volume (in orange) of a KUKA LBR iisy robot. The exploration of the free workspace has been performed through hand guidance with a cube-shaped exploration tool.
Our main contribution is a novel data-driven approach to modeling a constrained robot cell using exploratory robot motions, relying solely on the integrated joint encoders and eliminating sim-to-real gaps. The resulting swept volume is used to compute a (conservative) 3D mesh representation of the obstacles and unexplored areas, which is finally employed for automatic collision-free motion planning and control [6, 7]. Our approach is fast, cost-effective, and straightforward, supporting flexible production lines with confined spaces in modern Industry 4.0 scenarios. We validate the effectiveness of our approach in a pick-and-place scenario with the KUKA LBR iisy, demonstrating the user-friendly interface that does not require any advanced technical knowledge in robotics. Note that the approach is robot-agnostic, making it applicable to both collaborative and industrial robots.
@@
In the following, Sec. II presents the state of the art related to environment modeling and swept volume computation. Sec. III describes our approach to obtaining a representation of the free space within the constrained robot cell. Experimental results are reported in Sec. IV, while Sec. V concludes.
@@ III Free Workspace Exploration
This section describes the entire pipeline for modeling the constraints imposed by static obstacles in the robot cell environment. It is based on the robot’s exploratory motions without the use of any additional sensor. Our pipeline consists of four main steps (see Fig. 2), which will be detailed in the following subsections: 1) sweep through free space, 2) swept volume, 3) volume decimation, and 4) obstacle representation. The outcome is a mesh describing the boundary of the explored collision-free space. It can then be utilized within established collision-free trajectory planning and control methods to generate safe robot motions. Steps 1 – 3 can optionally be repeated to cover more space and improve the results.
[…]
@@
Assumptions: The approach considers a robot with ${N}$ non-static rigid links connected by joints, potentially including tools temporarily attached. The configuration-dependent position and orientation of all links must be accessible. We assume that accurate meshes for all links and tools are provided. Furthermore, the robot cell must be static, i.e., there are no moving obstacles or humans entering the workspace. Notice that these are reasonable assumptions for many industrial production sites. Our approach is solely based on data gathered through exploratory motions within few minutes and does not require any external perception.
In this letter, we are focusing on fixed-base manipulators in confined spaces, but the approach applies identically to floating-base systems such as legged and mobile robots. To illustrate the pipeline steps throughout this section, we chose an ideal planar robot manipulator with ${N}=3$ non-static links as a running example (Fig. 3(a)). Sec. IV reports results obtained with a KUKA LBR iisy collaborative robot.
@@ III-A Sweep through Free Space
The novel idea is to utilize the robot swept volume (SV) resulting from exploratory motions to identify the collision-free space within the constrained robot cell. During this exploratory phase, the robot’s volume may be intentionally modified, i.e., by mounting an additional rigid body on the robot flange. Increasing the robot’s overall shape allows for exploring more space in the same amount of time or the same space in less time without changing the robot’s velocity. A suitable modification also supports industrial scenarios requiring robot tool changes to satisfy the desired task scenario.
@@ III-A2 Exploration Tool
The end-effector tool(s) required for the actual task scenario can be used in the exploratory phase. However, the CAD models of such tools are typically very detailed, even when simplified, thus leading to high computation times and, possibly, to numerical instabilities during the subsequent steps of our pipeline. Therefore, in the exploratory phase, the robot’s shape may be modified by mounting an exploration tool on the flange, which increases the robot’s overall volume and speeds up the exploration process. Such an exploration tool can be customized for the exploratory motion phase. For its design, we recommend simple geometric shapes made of lightweight material without any electronics involved. Furthermore, choosing a form that encloses the tools (e.g., gripper or pneumatic suction cup) used later to achieve the actual task may be beneficial. The dimensions of the exploration tool may also be determined by considering the specific robot cell setup. If, for example, narrower gaps are of interest, the dimensions of the exploration tool should be designed accordingly. Moreover, the exploratory phase may also involve multiple exploration tools. Note that modern tool change systems allow users to quickly and flexibly switch between different tools. In the future, robot manufacturers might deliver new manipulators together with a set of such exploration tools that are cheap to produce.
@@ III-B Swept Volume
Given the joint trajectories ${\bm{q}\left(t\right)}$ recorded in the previous step, the corresponding $i$-th robot link pose ${\bm{T}_{i}\left(t\right)}$ at time $t$ is obtained by utilizing the robot kinematic model. This allows computing the swept volume ${V_{i}}$ of the $i$-th non-static robot link using (1). The computation can be performed for all ${N}$ non-static links in parallel improving computational efficiency [^1] Note that the base link of a fixed-base manipulator does not move, and, hence, its swept volume is identical to its body shape.. Note that the SV computation is not negatively affected by clustered or heterogeneously distributed data (e.g., same or similar joint configurations recorded in the exploration phase).
@@
Figs. 3(b) – 3(d) show the highly non-convex swept volumes associated with the three links of a planar robot, representing altogether the collision-free space discovered during the exploratory phase. No exploration tool has been used.
@@ III-D Obstacle Representation
${V_{O}}=\left(\left(\left({V_{BV}}~{}\backslash~{}{V_{1}}\right)~{}\backslash~{}{V_{2}}\right)~{}\backslash~{}\ldots~{}\right)~{}\backslash~{}{V_{{N}}},$ (2)
where ${V_{BV}}$ denotes a bounding volume covering the entire robot workspace. An intuitive choice for the bounding volume in the case of a robot manipulator could be a sphere or cube parameterized according to the robot size. The sequential operations in (2) are performed using the Boolean difference operator. Initially defined in the Boolean algebra, Boolean operations have extended to computer graphics and 3D modeling. The Boolean difference of two solid models $\mathcal{A},\mathcal{B}\subset\mathbb{R}^{3}$ is defined as $\mathcal{A}\,\backslash\,\mathcal{B}\coloneq\left\{\bm{x}\in\mathbb{R}^{3}\mid\bm{x}\in\mathcal{A}~{}\mathrm{and}~{}\bm{x}\notin\mathcal{B}\right\}.$ Given ${N}$ swept volumes associated with the non-static robot links, the Boolean difference operation in (2) is performed ${N}$ times [^2] The static base link of a fixed-base manipulator is not considered in (2).. The resulting volume ${V_{O}}$ corresponds to the bounding volume ${V_{BV}}$ without a portion of the inner part. Hence, ${V_{O}}$ can be interpreted as a conservative mesh representation including all obstacles and unexplored areas in the confined workspace. The robot is guaranteed to be collision-free as long as it does not penetrate ${V_{O}}$.
@@ IV Experiment
We validate the proposed approach within a pick-and-place scenario, confirming its effectiveness.
@@ IV-A Software Libraries
We selected the visualization toolkit [15] for the decimation algorithm. The input is a triangle mesh, and the parameters defining the decimation process (e.g., the target percentage of the triangle reduction, the maximum allowed error, and whether the mesh topology shall be preserved). It returns a triangle mesh.
Finally, we chose the software library libigl for the Boolean difference operation, which is based on [16] and cgal library666”CGAL, Computational Geometry Algorithms Library,” https://www.cgal.org/.. The algorithm inputs are two triangle meshes, provided as a collection of vertices and faces.
@@ IV-B Hardware Setup
The experimental platform is a KUKA LBR iisy 3 R760 with ${N}=6$ non-static links. It is equipped with a SCHUNK change system (FWA series), enabling a quick switch between the exploration tool and the gripper required for the task. The latter is a SCHUNK gripper (GEI FWA-50 series) with parallel 3D-printed fingertips. Overall, it approximately measures ${\left(7\times 7\times 12\right)\,\text{cm}}$. We designed a cube-shaped exploration tool to speed up the exploration phase and avoid high computation times or numerical instabilities. Fig. 6 shows the exploration tool, a lightweight cardboard box of ${\left(20\times 20\times 20\right)\,\text{cm}}$, which encloses the gripper entirely. Fig. 4 highlights the advantages of using this tool by comparing the robot SV mesh for a simple trajectory in three different scenarios: the robot as it is, with a parallel gripper, and with the exploration tool. The SV mesh generated with the gripper shows imperfections such as holes and irregularities. In contrast, the use of the exploration tool results in a smoother and more expansive SV. This underlines our secondary contribution.
Utilizing Robot Operating System (ROS) 2 to exchange data, all computations are performed on a laptop system with an Intel Core i7-12800H (2.4 GHz) CPU and 32 GB of RAM.
@@ IV-D Pick-and-Place Task
The robot must pick the orange blocks from the ramp and place them into the glasses. Each time the block at the bottom of the ramp is picked, the remaining ones slide down. This process is repeated six times until the ramp is empty. The first three blocks go into different glasses, and the next three follow the same order. In the end, each glass contains two blocks.
@@ IV-D4 Motion Planning and Control
Thanks to the change system, once the exploratory phase has been performed, the gripper quickly replaces the exploration tool. The pick-and-place task is modeled as a simple finite state machine, specifying the pick-and-place poses through the hand guidance mode. Moreover, since a collision is likely detected when the robot picks up the orange block, we record an additional pre-pick joint configuration $16$ cm above the ramp.
To generate optimal collision-free trajectories, we rely on the framework presented in [6] used with the flexible collision library (fcl) [18] for fast collision checks. Potential collisions are evaluated given the mesh of the robot, the gripper, and the environment represented as ${V_{O}}$.
@@ IV-E Execution Time Analysis
[…]
TABLE I: Execution times in seconds averaged over ten independent executions of the pipeline steps for a pick-and-place task with a kuka lbr iisy collaborative robot.
@@ IV-F The Role of the Volume Decimation
[…]
@@ V Conclusion
Future work involves integrating an autonomous exploration mode, where the robot changes its direction of motion upon contact detection. Furthermore, we would like to evaluate less accurate but faster AI-accelerated techniques similar to [11, 12] for visualizing the swept volume in real-time during an exploratory session (e.g., utilizing immersive augmented reality hardware). Additionally, the non-convex obstacle representation can be decomposed into multiple convex meshes using existing techniques, potentially accelerating collision checks for motion generation.

### v2

@@ I Introduction
Figure 1: Swept volume (in orange) of a KUKA LBR iisy robot (left: front view, right: back view). The exploration of the free workspace has been performed through hand guidance with a cube-shaped exploration tool.
Our main contribution is a novel data-driven approach to modeling a static constrained robot cell using exploratory robot motions, relying solely on the integrated joint encoders. Based on the collected data, we compute the robot swept volume, a non-convex hull represented as a 3D mesh that encapsulates the explored obstacle-free space, focusing on fixed obstacles. This mesh represents a conservative approximation of the accessible space. It can be utilized to verify if a particular robot configuration is safe. Hence, it can be integrated into state-of-the-art frameworks to enable automatic, collision-free motion planning and control [6, 7]. The assumption of a static environment is valid for many industrial applications, such as welding, painting, gluing, milling, deburring, and inspection. Our approach offers a compelling alternative to CAD-based modeling by eliminating sim-to-real gaps and reducing the need for complex modeling tasks. Moreover, it is fast, cost-effective, and straightforward – making it well-suited for supporting flexible production lines in modern Industry 4.0 scenarios. We validate the effectiveness of our approach in a pick-and-place scenario with the KUKA LBR iisy (see Fig. 1), highlighting the intuitive, user-friendly interface that does not require advanced technical knowledge in robotics. Note that the approach is robot-agnostic, making it applicable to both collaborative and industrial robots.
@@
A demonstration video of our work is provided as supplementary material and is available at https://youtu.be/p5DA41EtNUA.
In the following, Sec. II presents the state of the art related to environment modeling and swept volume computation. Sec. III describes our approach to obtaining a representation of the obstacle-free space within the constrained robot cell. It also addresses the limitations of our method and potential resolution strategies to overcome them. Experimental results are reported in Sec. IV, while Sec. V concludes.
@@ III Free Workspace Exploration
This section describes the entire pipeline for modeling the constraints imposed by static obstacles in the robot cell environment. The proposed method is entirely based on data acquired through exploratory motions, typically lasting only a few minutes, and does not require any additional external sensor. Our pipeline consists of four main steps (see Fig. 2), which will be detailed in the following subsections: 1) sweep through free space, 2) swept volume, 3) volume decimation, and 4) obstacle representation. The outcome is a mesh describing the boundary of the explored collision-free space. Steps 1 – 3 can optionally be repeated to explore additional free space and enhance the resulting model.
It is important to note that, similarly to CAD-based environment modeling, our approach does not address self-collisions or collision-free motion planning. Instead, it provides an environment model of the robot cell that can be integrated into established trajectory planning and control frameworks to generate safe robot motions.
Sweep through free space (Sec. III-A) Swept volume (Sec. III-B) Volume decimation (Sec. III-C) Obstacle representation (Sec. III-D) $\bm{q}(t)$$V_{i}$$\widetilde{V}_{i}$$V_{O}$
@@
Assumptions. The approach involves a fixed-based robot with N{N} non-static rigid links interconnected by joints, and is based on the following assumptions:
  1. A1.
    The robot cell is static, i.e., there are no moving obstacles or humans within the workspace;
  2. A2.
    Joint encoders are calibrated, and the configuration-dependent position and orientation of all links are accessible through precise forward kinematics;
  3. A3.
    Accurate meshes for all robot links and tools are available.
The pipeline steps will be illustrated throughout this section through an ideal planar robot manipulator with N=3{N}=3 non-static links and a KUKA LBR iisy collaborative robot.
@@ III-A Sweep through Free Space
The novel idea is to utilize the robot swept volume (SV) resulting from exploratory motions to identify the collision-free space within the constrained robot cell. During the exploratory phase, the robot’s volume may be intentionally modified, i.e., by mounting an additional rigid body on the robot flange. Increasing the robot’s overall shape allows for exploring more space in the same amount of time or the same space in less time without changing the robot’s velocity. The modification of the robot volume also supports industrial scenarios requiring robot tool changes to satisfy the desired task scenario.
@@ III-A2 Exploration Tool
The end-effector tool(s) required for the actual task scenario can be used in the exploratory phase. However, the CAD models of such tools are typically very detailed, even when simplified, thus leading to high computation times and, possibly, to numerical instabilities during the subsequent steps of our pipeline. Similar limitations can occur when utilizing a convex decomposition. Moreover, the exploration phase may become time-consuming in the case of small-size end-effector tools. Therefore, in the exploratory phase, the robot’s shape may be modified by mounting an exploration tool on the flange, which increases the robot’s overall volume and speeds up the exploration process. In this way, the exploration tool, rigidly connected to the robot flange, is considered part of the last link mesh. Such an exploration tool can be customized for the exploratory motion phase. For its design, we recommend simple geometric shapes made of lightweight material without any electronics involved. Furthermore, choosing a form that encloses the tools (e.g., gripper or pneumatic suction cup) used later to achieve the actual task may be beneficial. The dimensions of the exploration tool may also be determined by considering the specific robot cell setup. If, for example, narrower gaps are of interest, the dimensions of the exploration tool should be designed accordingly. Moreover, the exploratory phase may also involve multiple exploration tools. Note that modern tool change systems allow users to quickly and flexibly switch between different tools. In the future, robot manufacturers might deliver new manipulators together with a set of such exploration tools that are cheap to produce.
Utilizing a bounding box or a convex hull to represent the actual tool during the exploration phase constitutes a non-conservative approximation within our approach that can lead to severe problems. Both representations encompass spatial regions that are not part of the real tool – such as the gap between the two fingers of a parallel jaw gripper. Hence, utilizing such representations would erroneously label some regions as obstacle-free, even though they have not been traversed during the exploration phase (see Fig. 3). Consequently, this can lead to an incorrect representation of the obstacle-free space, which is unacceptable in safety-critical applications. The proposed exploration tool effectively addresses and overcomes this problem.
Figure 3: Example scenario where a gripper moves along a static metal bar, positioning its fingers on either side of the obstacle (left). Using the gripper’s convex hull (or bounding box) would incorrectly label the volume between the two gripper fingers as obstacle-free, despite being occupied (right).
@@ III-B Swept Volume
Given the joint trajectories $\bm{q}\left(t\right)$} recorded in the previous step, the corresponding $i$-th robot link pose $\bm{T}_{i}\left(t\right)$} at time $t$ is obtained by utilizing the robot kinematic model. This allows computing the swept volume $V_{i}$} of the $i$-th non-static robot link using (1). The computation can be performed for all N{N} non-static links – possibly including the exploration tool as part of the last link – in parallel improving computational efficiency [^1] Note that the base link of a fixed-base manipulator does not move, and, hence, its swept volume is identical to its body shape.. Note that the SV computation is not negatively affected by clustered or heterogeneously distributed data (e.g., same or similar joint configurations recorded in the exploration phase).
@@
Figs. 4(b) – 4(d) show the highly non-convex swept volumes associated with the three links of a planar robot, representing altogether the collision-free space discovered during the exploratory phase.
Fig. 5 illustrates examples of three-dimensional SVs of a KUKA LBR iisy manipulator performing three motions related to gluing and inspection tasks. The mesh details including vertices, faces, and volume [^2] The volume of the SV meshes has been computed with the 3d print toolbox of Blender., as well as the computation times, are reported in Table I.
@@ III-D Obstacle Representation
[…]
@@ III-F Discussion on Formal Guarantees
III-G Limitations
Our method is explicitly designed for deployment in static robot cell environments (Assumption A1). However, many real-world applications involve dynamic environments with moving obstacles and objects (e.g., on a conveyor), which can lead to unintended robot collisions. To automate processes in these environments, we suggest integrating the environment model generated by our approach with real-time data from additional sensors to detect and track dynamic obstacles and/or objects. On the other hand, it is worth noticing that the assumptions A2 and A3 generally hold in industrial applications.
Currently, our software implementation is not capable of real-time processing. Therefore, it is not possible to provide visual feedback of the growing robot swept volume during the exploration phase, though such feedback would be beneficial for the operator. We expect that future advancements in swept volume computation will help address this limitation.
@@ IV Experiment
We validate the proposed pipeline within a pick-and-place scenario, confirming its effectiveness.
TABLE I: Swept volume mesh details and computation time (ct) for a kuka lbr iisy 3 performing three different motions.
Motion Vertices Faces Volume CT
rect. path $19399$ $38810$ $0.43\,\text{m}^{3}$} $24.53\,\text{s}$}
semi-circ. path $13093$ $26182$ $0.29\,\text{m}^{3}$} $14.35\,\text{s}$}
shelf expl. $26724$ $53502$ $0.69\,\text{m}^{3}$} $57.90\,\text{s}$}
@@ IV-A Software Libraries
We selected the visualization toolkit [15] for the decimation algorithm. The input is a triangle mesh, and the parameters defining the decimation process (e.g., the target percentage of the triangle reduction, the maximum allowed error, and whether the mesh topology should be preserved). It returns a triangle mesh.
Finally, we used the software library libigl for the Boolean difference operation, which is based on [16] and cgal library777”CGAL, Computational Geometry Algorithms Library,” https://www.cgal.org/.. The algorithm inputs are two triangle meshes, provided as a collection of vertices and faces.
Figure 5: Swept volumes of a KUKA LBR iisy 3 for three different motions: a rectangular (left) and semi-circular (middle) path for a gluing task, and the exploration of a shelf (right) for an inspection task.
@@ IV-B Hardware Setup
The experimental platform is a KUKA LBR iisy 3 R760 with N=6{N}=6 non-static links. The robot is equipped with a SCHUNK change system (FWA series), enabling a quick switch between the exploration tool and the gripper required for the task. The latter is a SCHUNK gripper (GEI FWA-50 series) with parallel 3D-printed fingertips. Overall, it approximately measures $\left(7\times 7\times 12\right)\,\text{cm}$}. We designed a cube-shaped exploration tool to speed up the exploration phase and avoid high computation times or numerical instabilities. Fig. 9 shows the exploration tool, a lightweight cardboard box of $\left(20\times 20\times 20\right)\,\text{cm}$}, which encloses the gripper entirely. Fig. 6 highlights the advantages of using this tool by comparing the robot SV mesh for a planar trajectory in three different scenarios: the robot as it is, with a parallel gripper, and with the exploration tool. The SV mesh generated with the gripper shows imperfections such as holes and irregularities. In contrast, the use of the exploration tool results in a smoother and more expansive SV. This underlines our secondary contribution. Fig. 7 displays additional SV examples involving three distinct exploration tool geometries within a simple trajectory. Utilizing Robot Operating System (ROS) 2 to exchange data, all computations are performed on a laptop system with an Intel Core i7-12800H (2.4 GHz) CPU and 32 GB of RAM.
Figure 7: Swept volumes of a KUKA LBR iisy 3 executing a simple trajectory with three distinct exploration tool geometries: a cube (left), a cylinder (middle), and a triangular prism (right).
@@ IV-D Pick-and-Place Task
The robot must pick the orange blocks from the ramp and place them into the glasses. Whenever the block at the bottom of the ramp is removed, the remaining blocks shift downward, preserving a constant picking location. This process is repeated six times until the ramp is empty. The first three blocks go into different glasses, and the next three follow the same order. In the end, each glass contains two blocks.
@@ IV-D4 Motion Planning and Control
Thanks to the change system, once the exploratory phase has been performed, the gripper quickly replaces the exploration tool. The pick-and-place task is modeled as a simple finite state machine, specifying the pick-and-place poses through the hand guidance mode. Moreover, since a collision is likely detected when the robot picks up an orange block, we record an additional pre-pick joint configuration $16$ cm above the ramp.
Operators can flexibly select from a range of motion planning frameworks, including QP-based methods, AI-based planners, global planners, or reactive strategies, depending on the demands of the particular application. In this experiment, we rely on the framework presented in [6] and the flexible collision library (fcl) [18] to generate optimal collision-free trajectories through fast collision checks. Potential collisions are evaluated given the mesh of the robot, gripper, and the environment represented as $V_{O}$}. The motion framework also considers robot self-collisions.
@@ IV-E Execution Time Analysis
A thorough exploration (performed by the first author) of the free space surrounding the robot took $162.27\,\text{s}$} ($\approx{2.70\,\text{min}}$). Table II reports the execution times of each pipeline step (see Sec. IV-D) averaged over ten independent pipeline executions. From a computational perspective, the most time-consuming operation is the computation of the robot link swept volumes, which takes $185.81\,\text{s}$} ($\approx{3.1\,\text{min}}$). On the other hand, the computation of the volume decimation is almost negligible, taking only a few seconds. Overall, the manual robot workspace exploration and the automatic generation of the environment model $V_{O}$} took $368.49\,\text{s}$} ($\approx{6.14\,\text{min}}$). Additionally, programming the particular application required approximately $1.5\,\text{min}$} to record the home robot configuration, set the pick-and-place poses, and generate a collision-free trajectory. In summary, the robot operator spent $7.64\,\text{min}$} to model the robot cell and setup the collision-free robot program.
TABLE II: Execution times in seconds averaged over ten independent executions of the pipeline steps for a pick-and-place task with a kuka lbr iisy collaborative robot.
@@ IV-F The Role of the Volume Decimation
In this ablation study, we evaluate the role of the volume decimation step in the pipeline from a computational load point of view. Therefore, we have repeated the execution time analysis conducted in Sec. IV-E without performing the volume decimation step. The obstacle representation step works directly with the SV meshes $V_{i}$} produced by the previous step. The execution times to obtain the obstacle representation increase from $18.25\,\text{s}$} to $37.40\,\text{s}$}. Hence, the optional and almost costless execution of $2.16\,\text{s}$} for the volume decimation step implies a reduction of $16.99\,\text{s}$} ($\approx 4.41\,\%$) on the whole execution time. It is worth noticing that computational optimization may be significantly higher in the case of longer explorations or a more complex SV mesh.
@@ V Conclusion
Future work involves integrating an autonomous exploration mode, where the robot changes its direction of motion upon contact detection. Furthermore, we would like to evaluate less accurate but faster AI-accelerated techniques similar to [11, 12] for visualizing the swept volume in real-time during an exploratory session (e.g., utilizing immersive augmented reality hardware). Finally, a user study demonstrating the programming efficiency compared to conventional methods could offer valuable insights to the community.